the length of fence required to enclose a rectangular field is 300 meters. What are the dimensions of the field if it is known that hte difference between its lenght and width is 50 meters
Answer by uday1435(57)

Perimeter p = 2(l+w) = 300, Dividing by 2 we have l + w = 150 .......Eq.1
and l – w = 50 ..................................................... Eq.2
Adding Eq.1 and Eq.2 we get
2l = 200. , Dividing both sides by 2 we get l = 100
Since l + w = 150 w = 150 – 100 = 50
The answer is Length = 100 meters and width = 50 meters

If you still require further clarifications please write to
udayakumar.t.r@gmail.com